OpenStudy (khally92):

Solve the system of the linear equation and interpret your solution geometrically 2x+y+2z-4=0 x-y-z-2=0 x+2y-6z-12=o. I have solved so many, am finding this one tricky, we use the elimination method to eliminate one of the variables..

OpenStudy (anonymous):

Look for the easies variable to eliminate, in this case is y, just sum the first two equations. The first equation remains always the same, the second is substituted by the new equatin obtained by doing the sum. You can do that because since both sides are equal, summing the same thing on both sides does not change the relationship between the variables. Do the same thing for the last equation with the first one, but remember to get rid of the same variable that you used before so that you can do the same thing again with the two new equations, eliminating the last variable in this last step.

OpenStudy (khally92):

sum the first two equations 3x+z-6=0 am on it.

OpenStudy (khally92):

u mean i should subtract equation 1-3

OpenStudy (khally92):

equation 1-3 x-y+8z+8=0. nothing was eliminated.. if thats what you asked me to do?

OpenStudy (anonymous):

When you dont have anything that is eliminated automaticaly, you multiply the equations to get something you can eliminate. Like that: EQ3-2*EQ1 as you can see, the y will be aliminated because you have 2y on the EQ3 and y on EQ1. When you have something even more complicated (wich is not the case), you might need to multiply both equations to eliminate something.

OpenStudy (khally92):

yeah i add up the two equation with eliminated y. and i finally arrived at an equation ehich i solved for z=-10/9. i need more tip this time around. thanks. i have to submit this assignment 2morrow, thanks